OpenStudy (anonymous):

Write an equation in point-slope form for the line that passes through the pairs of points (5,1) and (-3,4)

OpenStudy (anonymous):

ok so do you know what point-slope form is?

OpenStudy (anonymous):

not really! slope intercept form is slope = y2-y1 over x2-x1 I think

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so you are going to find the slope

OpenStudy (anonymous):

so m=(4-1)/(-3-5) right?

OpenStudy (anonymous):

so m=3/-8 ?

OpenStudy (anonymous):

point-slope form is y-y1=m(x-x1)

OpenStudy (anonymous):

so you just choose a point whether it be (5,1) or (-3,4) and take the slope and plug it in to the formula

OpenStudy (anonymous):

How about y-1=-3/8(x-5) for slope form. Does that sound right?

OpenStudy (anonymous):

true dat :)

OpenStudy (anonymous):

I understand that - now how do you put those numbers in slope-intercept form?After that how do you put it in standard form?

OpenStudy (anonymous):

well slope-intercept form is y=mx+b soo first you would distribute the -3/8 so it would be y-1=(-3/8)x+(15/8) right?